# LT-Soups: Bridging Head and Tail Classes via Subsampled Model Soups

### v1

@@ Abstract
Real-world datasets typically exhibit long-tailed (LT) distributions, where a few head classes dominate and many tail classes are severely underrepresented. While recent work shows that parameter-efficient fine-tuning (PEFT) methods like LoRA and AdaptFormer preserve tail-class performance on foundation models such as CLIP, we find that they do so at the cost of head-class accuracy. We identify the head-tail ratio, the proportion of head to tail classes, as a crucial but overlooked factor influencing this trade-off. Through controlled experiments on CIFAR100 with varying imbalance ratio ($\rho$) and head-tail ratio ($\eta$), we show that PEFT excels in tail-heavy scenarios but degrades in more balanced and head-heavy distributions. To overcome these limitations, we propose LT-Soups, a two-stage model soups framework designed to generalize across diverse LT regimes. In the first stage, LT-Soups averages models fine-tuned on balanced subsets to reduce head-class bias; in the second, it fine-tunes only the classifier on the full dataset to restore head-class accuracy. Experiments across six benchmark datasets show that LT-Soups achieves superior trade-offs compared to both PEFT and traditional model soups across a wide range of imbalance regimes.
@@ 1 Introduction
In machine learning, balanced class distributions are often assumed in both theory and practice deng2009imagenet; 7968387; krizhevsky2009learning. However, real-world datasets frequently deviate from this assumption, exhibiting severe class imbalance where a few head classes dominate and tail classes remain significantly underrepresented van2018inaturalist; Holste2022LongTailedCO; liu2019large. This imbalance poses a fundamental challenge: models must learn effectively from limited tail-class data while preserving overall robustness chen2024survey.
Recent advances in vision-language foundation models, particularly CLIP radford2021learning, have introduced promising tools for addressing class imbalance. Trained on large-scale, diverse datasets, CLIP demonstrates strong robustness to distributional shifts and has become a popular backbone for long-tailed recognition wen2024what; wang2024exploring; ma2021simple; tian2022vl; long2022retrieval. Building on this, shi2024long achieve state-of-the-art results by applying parameter-efficient fine-tuning (PEFT) methods such as Low-Rank Adaptation (LoRA) hu2021lora and AdaptFormer chen2022adaptformer, in combination with logit adjustment (LA) loss menon2020long; ren2020balanced, which incorporates class priors by adding a class-dependent offset to the logits. While this PEFT-based approach improves overall and tail-class accuracy, they observe that it still underperforms full fine-tuning in certain regimes.
@@
These observations motivate a deeper investigation into when and why PEFT is effective. To this end, we construct a controllable long-tailed benchmark based on CIFAR100 that allows systematic variation in both the sample counts across classes, quantified by the imbalance ratio ($\rho$), and the number of classes above or below a sample threshold, defined as the head-tail ratio ($\eta$). This setup enables a more fine-grained analysis of imbalance structures. Within this framework, we compare two full fine-tuning strategies against a PEFT baseline: full fine-tuning with logit adjustment (LA) and model soups wortsman2022model, which averages the weights of multiple LA-trained models initialized with different seeds and hyperparameters. Results from this benchmark confirm previous findings: on average, PEFT improves overall accuracy (80.8 vs. 81.2) and tail-class accuracy (76.4 vs. 70.2) compared to full fine-tuning, but at the cost of degraded head-class performance (87.0 vs. 84.3). A detailed breakdown in Figure 1 shows that PEFT is especially effective in tail-heavy scenarios, where rare classes dominate ($\eta\ll 1$), but its performance declines as the class structure becomes more balanced or head-heavy, highlighting its limited robustness to shifts in class structure.
This highlights a key trade-off: PEFT helps prevent overfitting and supports tail classes, but lacks adaptability in more balanced and head-heavy settings. Conversely, full fine-tuning offers stronger adaptation but requires careful regularization. Model soups offer a middle ground by averaging models trained with different seeds and hyperparameters lakshminarayanan2017simple; ko2023fair, but our experiments show that traditional soups, built from models trained on the same imbalanced dataset, still underperform in tail-heavy cases, as they tend to overemphasize head-class performance due to the dominance of high imbalance ratios.
@@ 2 Related Work
Class imbalance has traditionally been tackled through oversampling minority classes, undersampling majority classes, or applying reweighted loss functions such as focal loss or logit adjustment (LA) chawla2002smote; he2009learning; menon2020long. While effective in certain settings, these techniques often struggle under overparameterized models zhai2022understanding. Decoupled training frameworks further refine this process by separating representation learning and classifier training kang2019decoupling; zhang2021distribution, assuming biases lie primarily in the classifier layer. However, this assumption breaks down when adapting foundation models, as full fine-tuning can lead to catastrophic forgetting and degraded generalization for both head and tail classes mukhoti2023fine; shi2024long.
Ensemble-based methods address class imbalance by combining experts trained on diverse data distributions cai2021ace; tao2023local. Examples include BBN zhou2020bbn and RIDE wang2020long, which use architectural branching or dynamic routing, and LFME xiang2020learning, which employs group-wise distillation. Mixture-of-Experts approaches such as SADE zhang2022self, Mdcs zhao2023mdcs, and DirMixE yang2024harnessing merge experts trained with different logit adjustments (e.g., uniform, long-tail, inverse long-tail). Unlike these methods that require all experts at inference, LT-Soups collapses multiple fine-tuned models into a single network via weight averaging, offering an inference-efficient alternative. While prior works rely on specialized architectures and heuristic expert definitions, our approach retains architectural simplicity by using parallel fine-tuning on controlled subsamples and model averaging to preserve both the generalization and efficiency of the foundation model.
CLIP and other vision-language models exhibit inherent robustness to class imbalance, largely due to the diversity of their pretraining data wen2024what. This robustness has been further extended through techniques such as prompt tuning dong2022lpt, retrieval-based augmentation long2022retrieval, and joint vision-language training paradigms ma2021simple; wang2024exploring. While these methods improve adaptation to long-tailed distributions, shi2024long show that PEFT combined with logit adjustment (LA) loss achieves state-of-the-art performance by selectively adapting CLIP’s pretrained features. However, they also observe that this comes at the cost of reduced head-class accuracy.
@@
Methods based on model merging, or weight averaging, has emerged as a practical strategy for reducing communication overhead in federated and distributed settings mcmahan2017communication; douillard2023diloco, improving robustness to distribution shifts wortsman2022model, and enhancing generalization through techniques like SWA and EMA izmailov2018averaging; tarvainen2017mean. Recent efforts also apply merging for continual learning and RLHF fine-tuning alexandrov-etal-2024-mitigating; rame2024warp. Yet, to our knowledge, model merging has not been explored for imbalanced classification.
@@ 3.1 Preliminaries
Given training data $\mathcal{D}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$, where $\boldsymbol{x}_{i}$ denotes an input sample and $y_{i}\in\mathcal{C}$ is its corresponding class label from a set of $K=|\mathcal{C}|$ classes. Let $n_{j}$ denote the number of training samples for class $j$, and let the total number of training samples be $N=\sum_{j=1}^{K}n_{j}$. Without loss of generality, we assume that classes are sorted in decreasing order of frequency, i.e., if $i<j$, then $n_{i}\geq n_{j}$. In the imbalanced setting considered here, the most frequent class is significantly larger than the rarest one, such that $n_{1}\gg n_{K}$. To quantify this imbalance, we define the imbalance ratio as $\rho=n_{K}/n_{1}$. Following liu2019large, we categorize classes with more than 100 training samples ($n_{j}>100$) as head classes, and the rest as tail classes. [^1] For simplicity, we initially group all low-resource classes into a single tail category. In the experimental section, we further subdivide the tail into medium-shot and few-shot groups for consistency with prior work. Since we aim to achieve balanced performance across all classes, we report $\text{BalAcc}=\frac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}\text{Accuracy}(c)$ which equally weights performance on each class.
Our model is composed of two main components: a feature extractor and a classification head. For feature extraction, we adopt the CLIP vision encoder, implemented using a Vision Transformer (ViT) dosovitskiy2021an and parameterized by $\theta$. The representation for input $\boldsymbol{x}$ is given by $f_{I}(\boldsymbol{x};\theta)=\boldsymbol{z}$ where $\boldsymbol{z}$ is the extracted feature vector. The final class prediction is computed as $\hat{y}=\arg\max g(\boldsymbol{z};\omega)$ where $g$ denotes the prototypical classification head with parameters $\omega$ constructed from the CLIP text encoder. (see Appendix D for the full details).
Previous work suggests that training with standard Cross-Entropy loss with instance-balanced sampling often leads to head-class bias due to class imbalance cao2019learning; he2009learning. Logit Adjustment (LA) menon2020long addresses this by adding a class-dependent offset to the logits, thereby correcting for prior class frequencies as follows:
@@
where $g_{y}$ denotes the predictive logit of model on class $y$ and $\pi\in\Delta_{y}$ are estimates of the class priors $\mathbb{P}(y)$ based on the empirical class frequencies on the training data $D$. However, shi2024long observed that when fine-tuning CLIP models from pretrained weights using LA (referred to as Full-FT), the resulting class-conditional distributions can become inconsistent, particularly for tail classes. To mitigate this, they advocate for methods that preserve proximity to the pretrained initialization, leveraging PEFT strategies such as LoRA and AdaptFormer.
@@ 4 LT-Soups: Imbalanced Learning by Subsampled Model Averaging
The preceding toy experiment illustrates that the optimal method depends on the underlying class structure. In head-heavy distributions, full fine-tuning is particularly effective, as it adjusts all model parameters to capture the rich structure of frequent classes. In contrast, when tail classes dominate, PEFT approaches like LoRA and AdaptFormer (as used in LIFT shi2024long) perform better by preserving pretrained representations that generalize well under limited supervision. Motivated by this trade-off, our goal is to design a method that maintains balanced performance across both extremes, regardless of the imbalance pattern.
Ensemble methods such as model soups wortsman2022model (Figure 3(a)) have demonstrated effectiveness in improving both overall and minority-class performance lakshminarayanan2017simple; ko2023fair. However, as shown in the previous section, traditional soups, while outperforming single-model fine-tuning, remain suboptimal for the full spectrum of long-tailed distributions. This is especially true in tail-heavy scenarios, where they tend to overemphasize head-class performance due to the dominance of high imbalance ratios in the training data. We address this limitation with LT-Soups, a soups-based framework specifically designed for long-tailed distributions. Each model in LT-Soups is fine-tuned on a subset of the training data with a distinct, reduced imbalance ratio (Figure 3(b)). While such subsampling enhances tail-class learning chaudhuri2023does, it may omit valuable head-class information kim2020m2m. To balance this, we train models on subsets with gradually increasing imbalance levels, allowing each model to specialize in different regions of the long-tail spectrum. The final model is obtained by averaging these specialized models. To further recover any lost head-class information, we introduce a second stage where only the classifier head is fine-tuned on the full dataset using a class-balanced objective (e.g., LA loss). Pseudocode for LT-Soups is provided in Algorithm 1, and the remainder of this section details the full procedure.
@@
Subsampling is a common strategy for addressing class imbalance by reducing overrepresented head-class samples he2009learning; chaudhuri2023does. However, aggressive subsampling can discard useful head-class information, degrading overall performance kim2020m2m; chawla2002smote; shi2023re. To address this, we propose progressive subsampling, which incrementally increases the imbalance ratio across subsets. Each model is fine-tuned on a subset with a specific ratio, preserving tail-class data while managing head-class underutilization. We construct the subset sequence as $\left\{D_{\rho_{i}}\;\middle|\;\rho_{i}=2^{i},\;i\in\left\{0,1,2,\dots,\left\lceil\log_{2}(\rho)\right\rceil\right\}\right\}$ where $D_{\rho_{i}}$ is a dataset with imbalance ratio $\rho_{i}$. This yields a sparse sequence of subsets with exponentially increasing imbalance ratios, ensuring broad coverage while limiting the number of models in the soups procedure. In practice, we retain only the first $N$ subsets, as extremely high imbalance ratios tend to overly favor head classes and degrade tail-class performance.
The resulting $N$ models, each trained on a different subset, are merged using a recursive interpolation strategy. Given weights $\{\theta_{n}\}_{n=0}^{N}$, where $\theta_{0}$ is the pretrained model, LT-Soups recursively combines models via $\theta_{n}=(1-\lambda)\theta_{n}+\lambda\theta_{n-1}$. The interpolation coefficient $\lambda$ controls knowledge retention from previous stages. This procedure ensures (1) proximity to the pretrained model $\theta_{0}$, preserving CLIP’s zero-shot capabilities wortsman2022robust, and (2) smooth integration of head and tail class representations zhou2020bbn. As shown in Section 5.3, LT-Soups exhibits partial insensitivity to the choice of loss function, owing to the balance introduced by subsampling and model averaging. However, since each subset remains mildly imbalanced, albeit less so than the full training set, applying LA loss during fine-tuning further mitigates the effects of label distribution shifts.
@@
While weight averaging and subsampling help mitigate head-class dominance, fine-tuning large pretrained models can still lead to degradation in tail-class performance wen2024what. To address this, we maintain an exponential moving average (EMA) of model weights via $\theta_{ema}=(1-\mu)\cdot\theta_{ema}+\mu\cdot\theta,$ with a momentum coefficient $\mu=0.99$. EMA acts as a regularizer during training huang2017snapshot, promoting convergence to flatter minima izmailov2018averaging, which has been shown to enhance generalization, particularly for underrepresented classes.
Since subsampling reduces data per subset and introduces variance, we adopt a bootstrapping strategy inspired by bagging breiman1996bagging: for each subset, we train $M$ models on different bootstrap samples and uniformly average their weights. This stabilizes learning and yields more robust representations.
@@
To further recover head-class information lost during subsampling, we perform a final fine-tuning stage on the classifier head using the full training set. The backbone is frozen to preserve merged representations, and LA loss is applied to adjust decision boundaries based on label frequencies. This step improves head-class accuracy without harming tail-class performance, similar to calibration in two-stage LT methods kang2019decoupling.

### v2

@@ Abstract
Real-world datasets typically exhibit long-tailed (LT) distributions, where a few head classes dominate, and many tail classes are severely underrepresented. While recent work shows that parameter-efficient fine-tuning (PEFT) methods like LoRA and AdaptFormer preserve tail-class performance on foundation models such as CLIP, we find that they do so at the cost of head-class accuracy. We identify the head-tail ratio, the proportion of head to tail classes, as a crucial but overlooked factor influencing this trade-off. Through controlled experiments on CIFAR100 with varying imbalance ratio ($\rho$) and head-tail ratio ($\eta$), we show that PEFT excels in tail-heavy scenarios but degrades in more balanced and head-heavy distributions. To overcome these limitations, we propose LT-Soups, a two-stage model soups framework designed to generalize across diverse LT regimes. In the first stage, LT-Soups averages models fine-tuned on balanced subsets to reduce head-class bias; in the second, it fine-tunes only the classifier on the full dataset to restore head-class accuracy. Experiments across six benchmark datasets show that LT-Soups achieves superior trade-offs compared to both PEFT and traditional model soups across a wide range of imbalance regimes.
@@ 1 Introduction
In machine learning, balanced class distributions are often assumed in both theory and practice [12, 73, 28]. However, real-world datasets frequently deviate from this assumption, exhibiting severe class imbalance where a few head classes dominate and tail classes remain significantly underrepresented [56, 19, 34]. This imbalance poses a fundamental challenge: models must learn effectively from limited tail-class data while preserving overall robustness [9].
Recent advances in vision-language foundation models, particularly CLIP [43], have introduced promising tools for addressing class imbalance. Trained on large-scale, diverse datasets, CLIP demonstrates strong robustness to distributional shifts and has become a popular backbone for long-tailed recognition [59, 58, 37, 55, 35]. Building on this, Shi et al. [50] achieve state-of-the-art results by applying parameter-efficient fine-tuning (PEFT) methods such as Low-Rank Adaptation (LoRA) [20] and AdaptFormer [8], in combination with logit adjustment (LA) loss [40, 46], which incorporates class priors by adding a class-dependent offset to the logits. While this PEFT-based approach improves overall and tail-class accuracy, they observe that it still underperforms full fine-tuning in certain regimes.
@@
These observations motivate a deeper investigation into when and why PEFT is effective. To this end, we construct a controllable long-tailed benchmark based on CIFAR100 that allows systematic variation in both the sample counts across classes, quantified by the imbalance ratio ($\rho$), and the number of classes above or below a sample threshold, defined as the head-tail ratio ($\eta$). This setup enables a more fine-grained analysis of imbalance structures. Within this framework, we compare two full fine-tuning strategies against a PEFT baseline: full fine-tuning with logit adjustment (LA) and model soups [60], which averages the weights of multiple LA-trained models initialized with different seeds and hyperparameters. Results from this benchmark confirm previous findings: on average, PEFT improves overall accuracy (80.8 vs. 81.2) and tail-class accuracy (76.4 vs. 70.2) compared to full fine-tuning, but at the cost of degraded head-class performance (87.0 vs. 84.3). A detailed breakdown in Figure 1 shows that PEFT is especially effective in tail-heavy scenarios, where rare classes dominate ($\eta\ll 1$), but its performance declines as the class structure becomes more balanced or head-heavy, highlighting its limited robustness to shifts in class structure.
This highlights a key trade-off: PEFT helps prevent overfitting and supports tail classes, but lacks adaptability in more balanced and head-heavy settings. Conversely, full fine-tuning offers stronger adaptation but requires careful regularization. Model soups offer a middle ground by averaging models trained with different seeds and hyperparameters [29, 27], but our experiments show that traditional soups, built from models trained on the same imbalanced dataset, still underperform in tail-heavy cases, as they tend to overemphasize head-class performance due to the dominance of high imbalance ratios.
@@ 2 Related Work
Class imbalance has traditionally been tackled through oversampling minority classes, undersampling majority classes, or applying reweighted loss functions such as focal loss or logit adjustment (LA) [7, 17, 40]. While effective in certain settings, these techniques often struggle under overparameterized models [66]. Decoupled training frameworks further refine this process by separating representation learning and classifier training [25, 67], assuming biases lie primarily in the classifier layer. However, this assumption breaks down when adapting foundation models, as full fine-tuning can lead to catastrophic forgetting and degraded generalization for both head and tail classes [41, 50].
Ensemble-based methods address class imbalance by combining experts trained on diverse data distributions [4, 53]. Examples include BBN [74] and RIDE [57], which use architectural branching or dynamic routing, and LFME [62], which employs group-wise distillation. Mixture-of-Experts approaches such as SADE [68], Mdcs [70], and DirMixE [65] merge experts trained with different logit adjustments (e.g., uniform, long-tail, inverse long-tail). Unlike these methods that require all experts at inference, LT-Soups collapses multiple fine-tuned models into a single network via weight averaging, offering an inference-efficient alternative. While prior works rely on specialized architectures and heuristic expert definitions, our approach retains architectural simplicity by using parallel fine-tuning on controlled subsamples and model averaging to preserve both the generalization and efficiency of the foundation model.
CLIP and other vision-language models exhibit inherent robustness to class imbalance, largely due to the diversity of their pretraining data [59]. This robustness has been further extended through techniques such as prompt tuning [13], retrieval-based augmentation [35], and joint vision-language training paradigms [37, 58]. While these methods improve adaptation to long-tailed distributions, Shi et al. [50] show that PEFT combined with logit adjustment (LA) loss achieves state-of-the-art performance by selectively adapting CLIP’s pretrained features. However, they also observe that this comes at the cost of reduced head-class accuracy.
@@
Methods based on model merging, or weight averaging, has emerged as a practical strategy for reducing communication overhead in federated and distributed settings [39, 15], improving robustness to distribution shifts [60], and enhancing generalization through techniques like SWA and EMA [24, 54]. Recent efforts also apply merging for continual learning and RLHF fine-tuning [1, 44]. Yet, to our knowledge, model merging has not been explored for imbalanced classification.
@@ 3.1 Preliminaries
Given training data $\mathcal{D}=\{(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$, where $\boldsymbol{x}_{i}$ denotes an input sample and $y_{i}\in\mathcal{C}$ is its corresponding class label from a set of $K=|\mathcal{C}|$ classes. Let $n_{j}$ denote the number of training samples for class $j$, and let the total number of training samples be $N=\sum_{j=1}^{K}n_{j}$. Without loss of generality, we assume that classes are sorted in decreasing order of frequency, i.e., if $i<j$, then $n_{i}\geq n_{j}$. In the imbalanced setting considered here, the most frequent class is significantly larger than the rarest one, such that $n_{1}\gg n_{K}$. To quantify this imbalance, we define the imbalance ratio as $\rho=n_{K}/n_{1}$. Following [34], we categorize classes with more than 100 training samples ($n_{j}>100$) as head classes, and the rest as tail classes. [^1] For simplicity, we initially group all low-resource classes into a single tail category. In the experimental section, we further subdivide the tail into medium-shot and few-shot groups for consistency with prior work. Since we aim to achieve balanced performance across all classes, we report $\text{BalAcc}=\frac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}\text{Accuracy}(c)$ which equally weights performance on each class.
Our model is composed of two main components: a feature extractor and a classification head. For feature extraction, we adopt the CLIP vision encoder, implemented using a Vision Transformer (ViT) [14] and parameterized by $\theta$. The representation for input $\boldsymbol{x}$ is given by $f_{I}(\boldsymbol{x};\theta)=\boldsymbol{z}$ where $\boldsymbol{z}$ is the extracted feature vector. The final class prediction is computed as $\hat{y}=\arg\max g(\boldsymbol{z};\omega)$ where $g$ denotes the prototypical classification head with parameters $\omega$ constructed from the CLIP text encoder. (see Appendix D for the full details).
Previous work suggests that training with standard Cross-Entropy loss with instance-balanced sampling often leads to head-class bias due to class imbalance [5, 17]. Logit Adjustment (LA) [40] addresses this by adding a class-dependent offset to the logits, thereby correcting for prior class frequencies as follows:
@@
where $g_{y}$ denotes the predictive logit of model on class $y$ and $\pi\in\Delta_{y}$ are estimates of the class priors $\mathbb{P}(y)$ based on the empirical class frequencies on the training data $D$. However, Shi et al. [50] observed that when fine-tuning CLIP models from pretrained weights using LA (referred to as Full-FT), the resulting class-conditional distributions can become inconsistent, particularly for tail classes. To mitigate this, they advocate for methods that preserve proximity to the pretrained initialization, leveraging PEFT strategies such as LoRA and AdaptFormer.
@@ 4 LT-Soups: Imbalanced Learning by Subsampled Model Averaging
The preceding toy experiment illustrates that the optimal method depends on the underlying class structure. In head-heavy distributions, full fine-tuning is particularly effective, as it adjusts all model parameters to capture the rich structure of frequent classes. In contrast, when tail classes dominate, PEFT approaches like LoRA and AdaptFormer (as used in LIFT [50]) perform better by preserving pretrained representations that generalize well under limited supervision. Motivated by this trade-off, our goal is to design a method that maintains balanced performance across both extremes, regardless of the imbalance pattern.
Ensemble methods such as model soups [60] (Figure 3(a)) have demonstrated effectiveness in improving both overall and minority-class performance [29, 27]. However, as shown in the previous section, traditional soups, while outperforming single-model fine-tuning, remain suboptimal for the full spectrum of long-tailed distributions. This is especially true in tail-heavy scenarios, where they tend to overemphasize head-class performance due to the dominance of high imbalance ratios in the training data. We address this limitation with LT-Soups, a soups-based framework specifically designed for long-tailed distributions. Each model in LT-Soups is fine-tuned on a subset of the training data with a distinct, reduced imbalance ratio (Figure 3(b)). While such subsampling enhances tail-class learning [6], it may omit valuable head-class information [26]. To balance this, we train models on subsets with gradually increasing imbalance levels, allowing each model to specialize in different regions of the long-tail spectrum. The final model is obtained by averaging these specialized models. To further recover any lost head-class information, we introduce a second stage where only the classifier head is fine-tuned on the full dataset using a class-balanced objective (e.g., LA loss). Pseudocode for LT-Soups is provided in Algorithm 1, and the remainder of this section details the full procedure.
@@
Subsampling is a common strategy for addressing class imbalance by reducing overrepresented head-class samples [17, 6]. However, aggressive subsampling can discard useful head-class information, degrading overall performance [26, 7, 49]. To address this, we propose progressive subsampling, which incrementally increases the imbalance ratio across subsets. Each model is fine-tuned on a subset with a specific ratio, preserving tail-class data while managing head-class underutilization. We construct the subset sequence as $\left\{D_{\rho_{i}}\;\middle|\;\rho_{i}=2^{i},\;i\in\left\{0,1,2,\dots,\left\lceil\log_{2}(\rho)\right\rceil\right\}\right\}$ where $D_{\rho_{i}}$ is a dataset with imbalance ratio $\rho_{i}$. This yields a sparse sequence of subsets with exponentially increasing imbalance ratios, ensuring broad coverage while limiting the number of models in the soups procedure. In practice, we retain only the first $N$ subsets, as extremely high imbalance ratios tend to overly favor head classes and degrade tail-class performance.
The resulting $N$ models, each trained on a different subset, are merged using a recursive interpolation strategy. Given weights $\{\theta_{n}\}_{n=0}^{N}$, where $\theta_{0}$ is the pretrained model, LT-Soups recursively combines models via $\theta_{n}=(1-\lambda)\theta_{n}+\lambda\theta_{n-1}$. The interpolation coefficient $\lambda$ controls knowledge retention from previous stages. This procedure ensures (1) proximity to the pretrained model $\theta_{0}$, preserving CLIP’s zero-shot capabilities [61], and (2) smooth integration of head and tail class representations [74]. As shown in Section 5.3, LT-Soups exhibits partial insensitivity to the choice of loss function, owing to the balance introduced by subsampling and model averaging. However, since each subset remains mildly imbalanced, albeit less so than the full training set, applying LA loss during fine-tuning further mitigates the effects of label distribution shifts.
@@
While weight averaging and subsampling help mitigate head-class dominance, fine-tuning large pretrained models can still lead to degradation in tail-class performance [59]. To address this, we maintain an exponential moving average (EMA) of model weights via $\theta_{ema}=(1-\mu)\cdot\theta_{ema}+\mu\cdot\theta,$ with a momentum coefficient $\mu=0.99$. EMA acts as a regularizer during training [21], promoting convergence to flatter minima [24], which has been shown to enhance generalization, particularly for underrepresented classes.
Since subsampling reduces data per subset and introduces variance, we adopt a bootstrapping strategy inspired by bagging [2]: for each subset, we train $M$ models on different bootstrap samples and uniformly average their weights. This stabilizes learning and yields more robust representations.
@@
To further recover head-class information lost during subsampling, we perform a final fine-tuning stage on the classifier head using the full training set. The backbone is frozen to preserve merged representations, and LA loss is applied to adjust decision boundaries based on label frequencies. This step improves head-class accuracy without harming tail-class performance, similar to calibration in two-stage LT methods [25].